If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) + g(x).
Answer by jim_thompson5910(35256)

f(x) + g(x) = (3x + 5) + (2x - 9)

f(x) + g(x) = 3x + 5 + 2x - 9

f(x) + g(x) = (3x + 2x) + (5 - 9)

f(x) + g(x) = (3 + 2)x + (5 - 9)

f(x) + g(x) = (5)x + (-4)

f(x) + g(x) = 5x - 4